OpenStudy (anonymous):

how do you do the equation: a hemoglobin has a spherical shape with a diameter of 5.5 nm. What is the diameter of the hemoglobin molecule in meters? What is it in millimeters?

OpenStudy (anonymous):

A billion nanometers in one meter - 5.5 nm * 1 m / 1000000000 = 5.5 × 10^-9 meters A thousand millimeters in one meter - 5.5 × 10^-6 mm

OpenStudy (anonymous):

\[5.5 nm (10^{-9}m/1nm)=5.5(10^{-9})m\]